solve by addition method (elimination).
5x + 2y = -17
3x -2y = -7

Answer by amarjeeth123(569)

5x + 2y = -17............equation 1
3x -2y = -7..............equation 2
Adding equations 1 and 2 we get,
5x+3x+2y-2y=-17+(-7)
8x=-24
x=-3
Substituting in equation 2 we get,
-9-2y=-7
-2y=2
y=-1
x=-3 and y=-1 is the solution.